# A Coalition Game for On-demand Multi-modal 3D Automated Delivery System

### v1

@@ Abstract
In urban logistics, Unmanned Aerial Vehicles (UAVs) and Autonomous Delivery Robots (ADRs) present promising alternatives for user and eco-friendly delivery solutions, especially when it comes to on-demand delivery. We introduce a multi-modal autonomous delivery optimization framework as a coalition game for a fleet of UAVs and ADRs operating in two overlaying networks to address last-mile delivery in urban environments, including high-density areas, road-based routing, and real-world operational challenges. Establishing the 3D network, a centralized dispatch system is designed to assign optimal delivery modes and determine the routing of the vehicle. The problem is defined as multiple depot pickup and delivery with time windows constrained over operational restrictions, such as vehicle battery limitation, precedence time window, and building obstruction. Subsequently, the coalition game theory is applied to investigate cooperation structures among the modes to capture how strategic collaboration among vehicles can improve overall routing efficiency. To do so, a generalized reinforcement learning model is designed to evaluate the cost-sharing and allocation to different coalitions for which sub-additive property and non-empty core exist. Our methodology leverages an end-to-end deep multi-agent policy gradient method augmented by a novel spatio-temporal adjacency neighbourhood graph attention network and transformer architecture using a heterogeneous edge-enhanced attention model. Conducting several numerical experiments on last-mile delivery applications, the result from the case study in the city of Mississauga shows that despite the incorporation of an extensive network in the graph for two modes and a complex training structure, the model addresses realistic operational constraints and achieves high-quality solutions compared with the existing transformer-based and heuristics methods and can perform well on non-homogeneous data distribution, generalizes well on the different scale and configuration, and demonstrate a robust performance under stochastic scenarios subject to wind speed and direction.
@@ 1 Introduction
The recent increase in the demand for online shopping and e-commerce has created new challenges toward a resilient logistics supply, specifically in on-demand last-mile delivery in densely populated areas, making it challenging to keep up with the current means of transportation as viable solutions for timely delivery. Thus, reshaping urban logistics cooperatively and effectively can be done to cope with the expected level of service in such situations. In this regard, multi-modal delivery is identified as a promising solution, such as minivans and UAV collaboration (Li et al., 2022), trucks and UAVs (Colajanni et al., 2023), and multi-vehicle truck and ADR delivery (Ostermeier et al., 2023) have been indicated to reduce the utilization of distribution vehicles significantly, timing and operation costs rather than being used individually. Given the complexity of urban environments and dynamically changing conditions, UAV and ADR-based delivery systems are expected to be adopted in the near future (Li and Kunze, 2023). The adoption is further facilitated by the investment in the infrastructure of operational charging facilities, improvements in their technology like batteries and payload capacity, and the development of regulations to ensure safe operations. Moreover, these vehicles reflect superior performance due to their ability to circumvent uncertainties. For instance, Samouh et al. (2020) showed that the hybrid UAV-ADR delivery system offers a higher level of service in various demand and fleet size scenarios, leading to reliable delivery times, which is crucial for on-demand services. Hence, in this study, we leverage the benefit of utilizing UAVs and ADRs as cooperative modes to propose an on-demand multi-modal last-mile delivery framework.
In recent years, the last-mile delivery systems have drastically changed in terms of structure and operations due to the complex nature of the dynamic of urban areas and e-commerce. Besides, various opportunities exist to access certain products daily or hourly, like Amazon Prime service and Uber Eats. This situation stimulates practical solutions for urban multi-modal on-demand routing problems, especially promoting cooperation among these entities in highly populated environments to minimize operational costs and time delay penalties. To do so, we first define a last-mile delivery scope of this research and then implement a game theory model through different coalition groups to investigate the benefit of cooperative modes delivery.
In the context of on-demand delivery, this research considers the multi-depot, multi-vehicle pick-up and delivery problem with time windows (PDPTW) in food delivery. It shows how an NP-hard combinatorial optimization problem with two modes can be solved time-efficiently in the urban environment to account for how on-demand delivery is operated according to the vehicle graph network. We assume that the ADRs operate on the existing sidewalk/bike lane network. Similarly, a predefined virtual network is extended in the air (3rd dimension) for the operations of UAVs designed and managed by the municipality. We aim to minimize the delay in customers receiving their packages by a pickup and drop-off type of delivery through a pre-determined network. Two separate ground and aerial graph networks for ADRs and UAVs are set up, which can travel along the edges. Furthermore, practical assumptions such as weather fluctuation and limited battery and capacity of automated vehicles are considered to capture the realistic aspects of urban delivery.
This research leverages the reinforcement learning (RL) approach for optimization, where agents follow the Markov Decision Property (MDP). A key advantage of learning-based approaches, specifically RL, is their ability to generalize from trained models to solve VRP instances of varying sizes and complexities without redesigning the algorithm or manually adjusting parameters (Bogyrbayeva et al., ). Additionally, PDPTW for last-mile delivery is constrained by pairing and precedence relationships, hard pickup, and soft delivery time. Due to its NP-hard nature, it remains difficult for conventional methods, including exact and heuristic algorithms, to solve it optimally in a short computation time (Zong et al., 2022), in which these constraints limit the solution space. On the other hand, deep reinforcement learning (DRL) augmented with a transformer is being used to automatically learn the rules in traditional heuristic methods for solving routing problems, which produces appealing results with much faster computation (Li et al., 2021; Kool et al., 2018). Inspired by them, we aim to solve PDPTW for the various applications of on-demand services, such as food and medicine delivery, by proposing a centralized deep multi-agent reinforcement learning (MARL) approach for coordinating a multi-modal fleet conditioned on the urban environment using dynamic edge-enhanced graph attention network transformer architecture with heterogeneous attention following the design in Zhang et al. (2023b); Zhang et al. (2022).
Finally, the notion of the core coalition game is brought to this study first to evaluate the cooperating efficiency of two modes if not to operate individually, and second, to promote last-mile delivery companies to form a coalition and collaborate among their vehicles to distribute their operational cost effectively by saving it through fairly allocating the total cost to each coalition. The core coalition theory as a cooperative game theory examines how groups or coalitions form between agents so that no member is incentivized to break from the greater group. The core exists if cost shares in which every coalition member gains a fair share and no smaller group would profit more from acting individually (Chalkiadakis et al., 2022). In this regard, after training the RL model for multi-agent, we use test instances on the generalized model to obtain the core among different coalitions of ADRs and UAVs if the core exists. A coalition game can be formed if the characteristic function satisfies the sub-additivity condition in the grand coalition. A characteristic function evaluates the guaranteed values of all participants’ combined costs in a coalition. To ensure such a combination, theoretical properties of the core coalition will be demonstrated for a toy example and then validated through numerical experiments. Note that the core can sometimes be empty, meaning cooperation is not preferred. Thereby, it does not necessarily reduce the total cost or work separately as a more efficient approach.
@@
The rest of this paper is organized as follows: a review of relevant studies and an extensive review of current vehicle routing problems and core coalition game theory are provided in Section 2. Section 3 briefly discusses problem statements, network of operation, and mathematical formulation. Methodology, proposed architecture, coalitional game theory analysis, and core existence are described in Section 4. The application of our proposed framework on the case study and their experimental result are discussed in Section 5. Finally, Section 6 is dedicated to conclusions, final remarks, and future works.
@@ 2 Background
2.1 Vehicle Routing Problem
This section reviews the existing literature on traditional heuristic methods and learning-based approaches for routing problems. In addition, the coalition game theory application is reviewed.
Savelsbergh and Sol (1995) was first studied the pickup and delivery problem, and exact methods were proposed as solutions early on Ruland and Rodin (1997), concluding that the dynamic programming performed well on Single-vehicle Pickup and Delivery Problem (PDP) in small-scale instances. On the other hand, attempting to generate exact optimality suffers from heavy computation due to the exponential complexity. As a substitute, more researchers are turning to heuristics to generate approximate optimal solutions, which can significantly increase efficiency while incurring minor quality costs (Zong et al., 2022). For instance, Chu et al. (2021) considered how the food delivery platform can improve by using efficient mini-batching gradient and heuristic algorithms designed to solve the joint order assignment and routing problem of last-mile delivery service in on-demand delivery.
Given the dynamic nature of such an issue, assigning on-demand requests to the best available vehicle on a limited planning horizon is challenging. Liu (2019) comprehensively designed a dynamic rolling horizon for UAV food delivery. Though practical factors such as the orders’ location uncertainty, variable demand, carrying capacity, battery consumption, and battery swapping operation have been considered, the computational cost exponentially grows for larger networks.
Even if heuristic-based methods could produce near-optimal solutions in a more reasonable amount of time, they could be more robust in adapting to the new environment; when dealing with high dynamics, the real-time simulation of these methods still needs to be improved. Even though numerous heuristic-based methods are developed to compute near-optimal solutions, their improvement is not on a larger scale and problem size as they get more time-consuming, and finding efficient rules is still a challenge.
2.1.1 Machine Learning Methods
Lately, there has been a rising inclination to employ deep reinforcement learning for tackling Vehicle Routing Problems (VRPs), and a prevalent characteristic among these approaches is the utilization of a policy network featuring an encoder-decoder structure. Typically, the encoder maps the 2-dimensional locations of customers or nodes into a feature embedding to extract pertinent information from the data. Subsequently, the decoder addresses the problem through either construction or improvement methodologies (Ma et al., 2022). In the construction approach, the decoder initializes with an empty sequence and progressively selects nodes at each step to formulate a complete solution. Conversely, the decoder commences with an already complete initial solution in the improvement approach and continually chooses candidate nodes or heuristic operators for specific operations at each step. This iterative process aims to enhance the solution until the termination criterion is satisfied. A masking scheme is consistently applied to mask the visited or invalid nodes to ensure each customer is visited only once. These Deep Reinforcement Learning (DRL) models can generate higher-quality solutions when augmented with attention mechanisms or graph neural networks.

### v2

@@ Abstract
In urban logistics, Unmanned Aerial Vehicles (UAVs) and Autonomous Delivery Robots (ADRs) present promising alternatives for user and cooperative delivery solutions, especially when it comes to on-demand delivery. We introduce a multi-modal autonomous delivery optimization framework as a coalition game for a fleet of UAVs and ADRs operating in two overlaying networks to address last-mile delivery in urban environments, including high-density areas and time-critical applications. In particular, a centralized dispatch system is designed to assign optimal delivery modes and solve the vehicle routing problem. The problem is defined as multiple depot pickup and delivery with time windows constrained over operational restrictions, such as vehicle battery limitation, precedence time window, and building obstruction. Utilizing the coalition game theory, we investigate cooperation structures among the modes to capture how strategic collaboration can improve overall routing efficiency. To do so, a generalized reinforcement learning model is designed to evaluate the cost-sharing and allocation to different modes to learn the cooperative behaviour with respect to various realistic scenarios. Our methodology leverages an end-to-end deep multi-agent policy gradient method augmented by a novel spatio-temporal adjacency neighbourhood graph attention network using a heterogeneous edge-enhanced attention model and transformer architecture. Several numerical experiments on last-mile delivery applications have been conducted, showing the results from the case study in the city of Mississauga, which shows that despite the incorporation of an extensive network in the graph for two modes and a complex training structure, the model addresses realistic operational constraints and achieves high-quality solutions compared with the existing transformer-based and classical methods. It can perform well on non-homogeneous data distribution, generalizes well on different scales and configurations, and demonstrates a robust cooperative performance under stochastic scenarios across various tasks, which is effectively reflected by coalition analysis and cost allocation to signify the advantage of cooperation.
@@ 1 Introduction
In recent years, on-demand delivery has gained remarkable attention. The market trend for online shopping demonstrates that the worldwide e-commerce sales are expected to top $7$ trillion in 2025, growing $15\%$ annually in North America since 2022 (LVMTech, 2025). This could potentially cause unpredictable traffic congestion and delivery delays in urban networks as cities grow rapidly. To effectively cope with the resulting time delays, there is a growing push to reshape the on-demand urban delivery frameworks. For instance, in the meal delivery industry, on-demand delivery services like Uber Eats have been investigated by several studies to facilitate continuous operation, optimize routing efficiency and maintain customer satisfaction (Wang et al., 2023a; Chen et al., 2023; Mehra et al., 2023). Nevertheless, their model incorporates ground riders, which cannot completely mitigate the long-term urban-related delivery issues. In particular, urban infrastructure layout and limited road network capacity significantly impact the timely and efficient last-mile delivery. Therefore, scaling the ground vehicles and current means of transportation will add more load to traffic flow and cause increased congestion, pollution, and risk of accidents (Mohammad et al., 2023).
Other use cases include deliveries during medical emergencies and natural disasters (Shi et al., 2022), as well as demand-responsive and urgent deliveries in high-density areas (He et al., 2022). These cases underscore the need to leverage air mobility to alleviate congestion and shorten delivery times. In this regard, autonomous air delivery and unmanned aerial vehicles (UAVs) have become an attractive solution to reduce delivery time and cost, access remote and difficult-to-reach areas, and bypass city traffic and inaccessible roads (Archetti and Bertazzi, 2021). Furthermore, their flexibility allows congestion-free delivery without recurrent stops (Beliaev et al., 2023). A variety of UAV on-demand last-mile delivery in urban environments has been explored. For example, parcel and meal delivery considering city structure, although all paths are considered point-to-point direct lines and have been designed for problem-specific delivery (Liu, 2019; Elsayed and Mohamed, 2020). On the other hand, there exist studies of on-demand delivery in high-density areas which incorporated a conflict-free aerial network considering urban infrastructures rather than a straight line by building a layout-inspired and road-based network (Mohamed Salleh et al., 2018), corridor-based route network planning, (He et al., 2022), and optimal height delivery in multiple flight levels (Kim et al., 2024), though mostly simulated a small case network and real-world delivery scenarios like peak hours demands have not been considered. In this regard, a conflict-free, scalable aerial network independent of physical city structure is required for autonomous on-demand urban delivery.
Nevertheless, practical issues have been identified that degrade the UAVs’ performance and restrict them from being the most effective solution in urban delivery scenarios. Regardless of noise and privacy issues, air regulations like no-fly zones and operational layers, including limited capacity and battery, directly affect the level of service. Despite their fast and agile delivery, these shortcomings may not reflect UAVs’ point of strength in cluttered urban areas. For example, a limitation of physical landing pads exclusively for a UAV landing in case of very low-level urban airspace delivery (Doole et al., 2020), in addition to the excessive energy consumption required for several take-offs and landings in multiple short-range flights in nearby locations (Zhang et al., 2021). As an alternative, collaborative delivery is identified as one of the promising solutions to exploit the individual strengths of other vehicles to overcome UAVs’ practical challenges. Namely, truck-drone hybrid delivery has been proposed for no-fly zone areas for the limited payload (Jeong et al., 2019), and truck-and-robot multi-modal delivery (Ostermeier et al., 2023) is suggested to save more cost and reduce traffic congestion. However, when it comes to inner-city networks, where dense residential areas exist, trucks bring more congestion. Autonomous Delivery Robots (ADRs), on the other hand, are another delivery candidate. They are designed to operate at low speeds, e.g., pedestrian speed, safely share existing sidewalks and bike lanes with people and provide service for a limited network (Alfandari et al., 2022). As such, a more efficient multi-modal choice for such an environment, a hybrid UAV-ADR, is proposed in a simple showcase one-to-one matching, leading to a higher level of service Samouh et al. (2020). Hence, we leverage a collaborative UAV-ADR system following this study as a multi-modal delivery to address more realistic urban last-mile delivery scenarios.
In the context of on-demand delivery, the research problem aims to optimize the operational cost for the city-scale delivery of food or medical supplies, where packages from pickup locations, like restaurants and medical centers, are to be delivered by a third-party or online delivery service to customers’ delivery locations. Initially, the sequential decision-making optimization scheme is designed as a centralized controller to constitute the node visiting sequence of each vehicle that minimizes the vehicles’ travel time and customers’ waiting time during the peak hour demand in an urban network. Subsequently, a collaborative delivery is offered that benefits from the synergy and interaction of each mode in the system, first to distribute the workload between vehicles and secondly to investigate its impact on the cost savings of individual modes. We assume the ADRs operate on the existing sidewalk/bike lane network. Similarly, a predefined virtual network is extended in the air (3rd dimension) for the operations of UAVs managed by the municipality. Figure 1 represents a pickup and delivery environment for the UAVs’ operation network in the city of Mississauga. As a result, given the battery-based vehicles and realistic aspects of urban delivery, this problem can be classified as a collaborative electric capacitated multi-vehicle pick-up and delivery problem with time windows (CE-CPDPTW).
Figure 1: Aerial network for UAVs, including a restricted zone marked as a red arced region
It is noted that throughout this research, the term collaborative means that each mode works together and cooperates in the form of groups towards a common objective by collective decision-making rather than individual strategic choices. In the field of game theory, this is also called a coalition game and aims to model the players’ cooperation behaviour given heterogeneous agents and the complexity of the environment (Ahmad et al., 2023). Specifically, in urban last-mile delivery, collaborative delivery has gained increasing interest to reduce transportation costs and improve operational efficiency (Gansterer and Hartl, 2020; Zhang et al., 2022), especially to economically incentivize individuals’ cooperation by fair cost allocation to balance cost savings (Pingale et al., 2024). However, there is a gap in the collaborative UAV-ADR delivery literature, and customer service sharing in the CE-CPDPTW application of last-mile delivery has not been sufficiently studied, specifically in what real-world urban scenarios vehicle coalitions are beneficial and how their joint effort can be divided among each player. Therefore, solving CE-CPDPTW, we aim to obtain the sub-coalitions of two modes in which the cooperation gain is higher than the individual gain. While each participant can act independently with local observation, in this work, all players require full information sharing since they are managed by a central dispatcher.
In terms of the vehicle routing problem (VRP), heuristic methods were among the earliest approaches to solving the VRP variant. They are, however, frequently problem-specific and manually developed and cannot generalize or scale to other problem settings (Das et al., 2020; Alfandari et al., 2022; Gu et al., 2023). On the other hand, deep reinforcement learning (DRL) approaches have been presented as a powerful tool to model complex objectives given complicated constraints and uncertainty in last-mile delivery problems (James et al., 2019). Specifically, they can handle larger-scale problems with different constraints and outperform traditional and metaheuristic algorithms and generalize from trained models to solve VRP instances of varying sizes and configurations Li et al. (2022); Bogyrbayeva et al. (2022). Furthermore, the majority of the recent efforts on DRL-based VRP have employed encoder-decoder deep learning architectures to simulate sequential decision-making. For instance, transformers (Vaswani et al., 2017), for node sequence decoding (Kool et al., 2018) and graph attention networks (GAT) (Velickovic et al., 2017), for graph encoding (Lei et al., 2022), have exhibited promising alternatives in terms of computation time, generalization, and dealing with stochasticity due to capability of strong information embedding in graph structure using attention mechanisms. Inspired by them, we aim to solve CE-CPDPTW for the various applications of on-demand services, such as food and medicine delivery, by proposing a centralized deep multi-agent reinforcement learning (MARL) approach optimization. To model the sequential decision-making, we employ a heterogeneous, edge-enhanced graph attention network encoder based on temporal and spatial features and a transformer architecture decoder for a multi-modal fleet conditioned on the urban environment.
Finally, the notion of the core coalition game is brought to this study first to evaluate the cooperating efficiency of two modes if not to operate individually, and second, to promote single-mode last-mile delivery companies to horizontally collaborate upon sharing information or to form a coalition and benefit from both UAV and ADR to distribute the operational and customer service cost to each coalition. To the best of the author’s knowledge, this is the first study to propose a collaborative autonomous UAV-ADR fleet addressing the urban-related on-demand last-mile delivery applications, firstly, by introducing a dual GAT-transformer mechanism with priority-based and conflict-free mode assignment, followed by a coalitional evaluation and cost allocation mechanism to incentivize the agents to illustrate the advantage of working in coalitions across various case studies. To summarize, the main contributions of our work are listed as follows:
  • 1.
    We tackle the CE-CPDPTW problem for a multi-modal and multi-depot fleet with limited payload capacity and battery size operation for pickup and delivery problems with a time window to minimize the delay and travel time, which accounts for real-world urban delivery cases, such as time-critical and non-uniform distribution delivery, including weather uncertainty and urban layouts.
  • 2.
    We propose an end-to-end multi-agent reinforcement learning methodology with a dual edge-enhanced spatial-temporal-aware encoder incorporated with an extensive yet effective graph attention module to differentiate the multi-modal network. Then, we generate the visiting node sequence of the optimal travel tour by adaptive speed routing using a transformer decoder through a conflict resolution layer.
  • 3.
    Furthermore, a coalition game is defined to capture the interactive behaviour of the system to obtain potential coalitions to cooperate, if they exist, to suggest accommodating UAV-ADR combination and provide the managerial insights for online delivery services for performance attribution and regulatory trade-off of their fleet by evaluating the impact of the cost allocation mechanism of the savings on individual modes’ collaboration incentive.
@@
The rest of this paper is organized as follows: a review of relevant studies that have motivated our design choices from a methodological and collaborative perspective of urban last-mile delivery provided in Section 2. Section 3 briefly discusses problem statements, mathematical formulation, and coalitional game theory preliminaries. Methodology and detailed GAT-transformer architecture are described in Section 4. The application of our proposed framework through an extensive experimental result analysis, as well as cooperative game evaluation, is discussed in Section 5. Finally, Section 6 is dedicated to conclusions, final remarks, and future works.
@@ 2 Background
This section reviews the existing literature on the CE-CPDPTW routing problem, specifically when utilizing encoder and decoder DRL architectures. In addition, collaborative delivery strategies as well as the coalition game applications regarding the urban last-mile delivery are reviewed. The research on modelling the on-demand last-mile delivery applications and incorporating complex real-world constraints is well established. However, to the best of our knowledge, there are only a limited number of studies that focus on solving the CE-CPDPTW problem using two interactive modes of UAVs and ADRs, particularly addressing realistic urban delivery.
2.1 CE-CPDPTW Approaches
Chu et al. (2021) considered how food delivery platforms can solve the joint order assignment and routing problem of last-mile delivery service in on-demand delivery using efficient mini-batching gradient and simulated annealing algorithm. Moreover, Liu (2019) comprehensively designed a dynamic rolling horizon for UAV food delivery. Although practical factors such as the orders’ location uncertainty, variable demand, carrying capacity, battery consumption, and battery swapping operation have been considered, the computational cost exponentially grows for larger networks in both examples. Additionally, the CE-CPDPTW problem for last-mile delivery is constrained by pairing and precedence relationships, urban physical layout, and time window delivery. Due to its NP-hard nature, it remains difficult for conventional methods, including exact and heuristic algorithms, to solve it optimally in a short computation time (Zong et al., 2022), in which these constraints limit the solution space. In contrast, reinforcement learning (RL) approaches in combination with sequential deep learning models are proposed to automatically learn the rules in traditional heuristic methods for solving routing problems, which produces results with much faster computation.